your given the following national income model Y=C+I+G ,C=12+0.8Y ,I=100+0.1Y ,G=300



present the model to matrix formula.

1
Expert's answer
2022-02-04T09:20:05-0500

To transform to a matrix, the following guideline has to be adhered.

"Y=C+I+G\to""Y-C=I+G"

"C=a+bY\to" "C-bY=a"

"T=tY\to" "T-tY"

"Y=C+I+G\to Y=12+0.8Y+100+0.1Y+300"

"Y-0.9Y=412\to Y=4120\therefore I=112+0.9(4120)" "=512"

We then replace the values from the given equations where relevant, and placing a coefficient of zero to the missing constants in the guide equation.

"1\times Y-1\times C+0\times T=I+G"

"-0.8\times Y+1\times C + 0\times T=12"

"-t\times Y + 0\times C + 1\times T=0"

The matrix obtained from the variable coefficients become;

"A=" "\begin{bmatrix}
1 & -1 & 0 \\
-0.8 & 1 & 0 \\
-t & -0 & 1 \\
\end{bmatrix}" "X=" "\begin{bmatrix}
Y \\
C \\
T \\
\end{bmatrix}" "Z=" "\begin{bmatrix}
I+G \\
12 \\
0 \\
\end{bmatrix}" but I and G values are known.

"I+G=512+300=812"

so "Z=\begin{bmatrix}
812 \\
12 \\
0 \\
\end{bmatrix}"